You are going to purchase two types of computers for your business. If computer A costs $700 and computer B costs $500, then how many of each should you buy if you need 16 computers and you have $10,000 to spend?
Answer by priya(51)

First try to form equations based on the questions.
1. let X be the number of computer A and Y be the number the number of computers B
2.Now u know the total number of computers u r planning to buy..that is 16.
so the first euation would be X + Y = 16.
3. Now,try to form another equation with the help of the costs.
that is, 700 * X + 500 * Y = 10,000 ( THE TOTAL AMOUNT U R GOIN TO SPEND IS 10,000).So we've got 2 equations and now it is easy to solve the problem..
write the 2 equations:
EQ.1 700 * X + 500 * Y = 10,000
EQ.2 X + Y = 16
Now multiply the whole eq.2 by -500(always try to make the co-efficient of EITHER X OR Y in both equations to be the same with opposite sign,here we r tryin to make co-efficient of Y SAME ,BUT WITH OPPOSITE SIGN,THAT IS -500) ,and u'll get eq.2 as

new EQ.2 -500 * X -500 * Y = -8000 ,(this is done to solve the 2 equations and find X & Y)
NOW ADD EQ .1 AND THE NEW EQ.2.
We get ,
200 * X = 2000 (+500X -500X GET CANCELLED)
SO X = 2000/200
= 10.
WE KNOW THAT X + Y = 16,THEREFORE
10 + Y = 16 AND Y = 16- 10
Y= 6.
SO THE VALUES ARE NUMBER OF COMPUTER A IS 10 AND COMPUTER B IS 6.
WANT TO LEARN MORE !!VISIT MY SITE..